An ideal gas with a fixed number of molecules is m
Answer #1

P=2*105pa or 200kpa(1pa=0.001 Kpa)
V=6 m3
b=P*V
b=1200 kpam2
As b is contanst

so, when P'=4P
b=P'*V
V=b/P'=b/4P
=1200/4*200
=1.5 m3

V= volume
P= pressure
b= numeric constant